OpenStudy (anonymous):

solve for x (square root of 8) (square root of 2 - x) = 11

OpenStudy (anonymous):

\[\sqrt{8} * \sqrt{2-x} = 11\] \[\sqrt{8(2-x)} = 11\] \[\left( \sqrt{8(2-x)} \right)^2 = 11^2\] \[8(2-x) = 121\] Can you do it from here? Just distribute the 8 and then solve for x :)

OpenStudy (anonymous):

Uh, can you explain more? @AshleyAlley<3

OpenStudy (anonymous):

or could someone else give me the answer?

OpenStudy (luigi0210):

Distribute and solve